# vesselFM: A Foundation Model for Universal 3D Blood Vessel Segmentation

### v1

@@ 10 Additional Qualitative Zero-Shot Results
11 More Samples From $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$
We present additional samples from $\mathcal{D}_{\text{drand}}$ = $\{\mathcal{X}_{\text{drand}},\mathcal{M}_{\text{syn}}\}$ and $\mathcal{D}_{\text{flow}}$ = $\{\mathcal{X}_{\text{flow}},\mathcal{M}_{\text{syn}}\}$ in Fig. 11a and b, respectively. In conclusion, one can state that our domain randomized generative pipeline produces a wide variety of image-mask pairs with highly diverse fore- and background geometries and textures, while images sampled from our flow matching-based generative model $\mathcal{F}$ exhibit intensity patterns closely mimicking those of real images in $\mathcal{D}_{\text{real}}$.
12 Additional Flow Matching Trajectories
@@
13 More Masks $\mathcal{M}_{\text{syn}}$

### v2

@@ Appendix I Additional Qualitative Zero-Shot Results
Appendix J Additional Flow Matching Trajectories
@@
Appendix K Statistical Analysis of Results
We conducted a statistical analysis of the quantitative results reported in the main manuscript using paired t-tests. To this end, we compared vesselFM’s base configuration to the respective runner-up. We found all results to be statistically significant ($p$ $<$ 0.05), except for the Dice score in the ablation on class conditioning (Table 4, 4th row; $p$ = 0.081). The clDice score in the ablation on class conditioning, however, remains significant ($p$ = 0.036). Notably, values reported in Table 5 are statistically significant ($p$ = 0.019 for Dice; $p$ = 2.39 $\cdot$ $\text{10}^{-\text{5}}$ for clDice).
Appendix L Computational Resources
VesselFM comprises 31,418,977 parameters and requires minimal computational resources compared to other foundation models for 3D image segmentation. Processing a volume of shape $\text{128}^{\text{3}}$ takes 335.7 ms on a T4 GPU, 95.8 ms on a V100 GPU, 30.6 ms on an A100 GPU, and 3.3 s on an AMD Epyc 7702 CPU, with a VRAM consumption of approximately 4.21 GB.
Appendix M Comparison to Specialist Models
We further compare vesselFM to specialist models trained on individual blood vessel segmentation datasets, as shown in Table 7. As expected, vesselFM not only consistently outperforms state-of-the-art foundation models for medical image segmentation, but also specialist models, even when pre-trained on exactly the same imaging modality (CSD and SMILE-UHURA both contain MRA images).